OpenStudy (anonymous):

calculate ratio of distance travelled by body in 2nd and 4th second during free fall

OpenStudy (anonymous):

s= -half g * t^2..! now take ration of s1 and s2 and its equal to ratio of squares of the time intervals!

OpenStudy (unicorn):

attachement below

OpenStudy (anonymous):

distance travelled in 4th second=1/2 *g*(16-9)=7/2*g distance travelled in 2nd second=1/2*g*(4-1)=3/2*g ratio=3/7